IMACS progress report no. 52
B. Bigelow 12 August 2002
I) Project Status
Currently scheduled tasks: Status:
a) DSS drive design (Hare) IP
b) DSS drafting (except drive) (Cleary) Done
c) DSS assembly/testing (Hare, Kowal, Castillo) IP
d) DSS Grating Tilt Mech testing (Hare) IP
e) Calibration system mech. des. (Hare) on hold
f) CCD array assembly (Thompson/Burley) IP
g) Filter procurement (Sutin) Done
h) Grism procurement (Sutin) IP
i) Instrument electronics assembly (Bagish) on sched.
j) Control software design (Birk) on sched.
k) Instrument enclosures fab (Gunnels/Paragon) IP
l) Instrument utility wrap design (Gunnels) done
m) Utility wrap fabrication IP
Project events in July:
- First Light on 4 of 8 CCDs with long camera!
- 2/3 of enclosure panels delivered in July
Latest project images:
Final Assy: http://www.ociw.edu/instrumentation/imacs/devel/final/index.html
Dewar Assy: http://www.ociw.edu/instrumentation/imacs/devel/dewar/index.html
II) Optics (Sutin)
No report this month.
- Aspheres due from Tinsley in August.
- SO1 is "finished", and undergoing final testing.
- S08 is close behind.
- As-built update for short camera is scheduled for August
III) Electronics (Bagish)
- Assembled and successfully tested the new Altera hard stop logic, PI
encoder repeater, Microphone preamp, line driver and line receiver.
Final quantities of all PCB's have been ordered.
- Two MCS Rack bulkhead panels have been engraved and connectors installed.
- Assembly of 35 foot long cables has begun. Additional cable ordered and
- TCIP/IP to serial interface break-out schematic complete and parts ordered.
- Two RTD thermal sensors ordered for initial test.
- One-half of two-rack insulating panels have been cut and mortised to size.
- Bypassed thermostat on heat exchanger.
- Parts ordered and ready to test fluid flow of heat exchanger.
- Populate and test all remaining PCB's.
- Upgrade MCS Chassis 3 & 4 with new Altera hard-stop logic.
- Send MCS front panels out for final finish, anodizing and engraving.
- Upgrade MCS Chassis 1 & 2 with Altera hard-stop logic and new front panels.
- Test RTD thermal sensors w/ HP data acquisition unit.
- Assemble cables.
IV) Software (Birk/Clardy)
- Long Camera focus: We tested the Long Camera focus motors and were
able to focus an image of a round slit. Moving the focussed image in
X and Y turned out to greatly defocus the image. It does not seem
feasible to me to use LO5 for flexure control, but moving all 3
motors in parallel to correct for scale/focus changes works fine.
- Shutters: The shutter was connected to the DSP for the first time
and we were able to take pictures using the IMACS dewar.
- Gratings/Disperser-wheel: Lots of flexure tests were done this month
and major hardware changes are going on. I hope to continue with
software test on the disperser mechanisms in August.
- DSP/Science Array Software: For the first time I read CCD (MIKE)
data from the DSP. The full array mode works fine. There are onle
small changes necessary in the DSP code, eg. removal of prescan
pixels. The sub-raster- and shuffle-modes and reading more than one
CCD chip simultaneously, requires to update the DSP code. Greg and
I are trying to make those changes by email while he is in Canada.
- Graphical IMACS Status Display: I continue to add parts to the
graphical status display GUI. It now shows the hatch, all three
guiders, the slit-mask server, disperser server, both filter-servers
V) Detector Systems (Thomson/Burley)
First of all, the bad news -- one of the IMACS cryotigers had an
o-ring blow-out in the cooling head. Not much gas escaped, but it did
mean that one out of two IMACS cryotigers was out of action. It has
been sent back for servicing, along with the other cryotiger system
which needed the cooling gas exchanged from PT-13 to PT-30.
Also, we accidentally discovered a somewhat nasty thermal short in the
IMACS dewar. When the dewar was evacuated, the focus stage was pushed
towards the faceplate, and smashed the ball-bond covers on the SITe
detectors into the faceplate. This could have been a real disaster if
the ball-bond covers weren't there. There is a retaining nut on the
focus motor shaft that should prevent this happening, but the focus
motor assembly wasn't installed at the time.
And then, the good news -- both the Red MIKE + Blue MIKE ccd cameras
were installed and working on the spectrograph. And, roughly half the
IMACS mosaic was working with the long camera optics on that
spectrograph -- 4 chips + 1 cryotiger.
The single cryotiger was able to cool the detectors to about -90C,
which is much colder than might be expected.
1) putting the MIT/LL chip in blue MIKE
2) getting the fiber link to work (prereq for the 8-chip readout)
3) populating the rest of the chips in the IMACS array
VI) Mechanics (Bigelow/Hare)
1) Complete 2nd imaging mirror cell - done, parts in shop
2) Progress grism cell design - no progress
3) Designed and received spool which allows the long camera to
mount on the short camera end of the structure - flexure testing
4) Measured lens motions in cells - found to be within allowances
5) Completed final layout of cables and selected IGUS chain for
6) Wrote two papers for SPIE Astronomy 2002 meeting in August
1) Progress grism cells
2) On vacation 8/5 - 8/11
3) Short camera as-built update - tooling design
4) Order long camera baffles
5) SPIE meeting 8/19 - 8/29
* Continued optomechanical and electro-mechanical assembly and testing of the Principle
and Shack-Hartmann Guide Cameras.
* Assembly and testing of the Grating Tilt Mechanism. Investigating flexure issues.
* Review of the Disperser Server Drive Mechanism design complete. Revising solid
* Continue assembly, testing, and rework of all guide cameras as necessary.
* Continued testing of the Grating Tilt Mechanism.
* Fabrication DSS Drive Assembly.
* Assembly and testing of the Disperser Server System and Cable Wrap Subassembly.
* Fabrication of the second shutter.
* Resume design work of the Calibration Unit.
VII) Fabrication (Kowal)
1) Made new drive wheel on IMACS structure
2) Made new Drive Plates for Shack and Prin. Guiders
3) Made assorted Shutter parts.
4) Built test fixtures for Tilt Mechanism
5) Modified Tilt Mechanism Part for rigidity
6) Modified supports for mask server to clear enclosure panels
1) Shutter v2.0
2) New Imaging Mirror Cell
3) Pick up work
VIII) Other Business
Next meeting: 10:15, Wednesday 9/4/02